OpenStudy (anonymous):

WILL GIVE MEDAL AND FAN :) It takes a car 1.75 h to go from mile marker 10 to mile marker 115. What is the average speed of the car? A. 1.6 mph B. 60 mph C. 65 mph D. 183 mph

OpenStudy (anonymous):

B

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

time=1.75 hours distance=115-10 S=105miles S=vt v=S/t v=105/1.75 v=60mph

OpenStudy (anonymous):

Okay thank you.How do you give medals?

OpenStudy (anonymous):

I dont really need a medal its fine (:

OpenStudy (anonymous):

No its not fine! you helped me.

OpenStudy (anonymous):

Thanks for the medal (: